# VLA-Mark: A cross modal watermark for large vision-language alignment models

### v1

@@ Abstract
Vision-language models demand watermarking solutions that protect intellectual property without compromising multimodal coherence. Existing text watermarking methods disrupt visual-textual alignment through biased token selection and static strategies, leaving semantic-critical concepts vulnerable. We propose VLA-Mark, a vision-aligned framework that embeds detectable watermarks while preserving semantic fidelity through cross-modal coordination. Our approach integrates multiscale visual-textual alignment metrics, combining localized patch affinity, global semantic coherence, and contextual attention patterns, to guide watermark injection without model retraining. An entropy-sensitive mechanism dynamically balances watermark strength and semantic preservation, prioritizing visual grounding during low-uncertainty generation phases. Experiments show 7.4% lower PPL and 26.6% higher BLEU than conventional methods, with near-perfect detection (98.8% AUC). The framework demonstrates 96.1% attack resilience against attacks such as paraphrasing and synonym substitution, while maintaining text-visual consistency, establishing new standards for quality-preserving multimodal watermarking  [^1] Code will be available at https://github.com/shiningwhite-cmd/VLA-mark.
@@ 1 Introduction
The emergence of vision-language aligned multimodal large models (VLAMMs) has fundamentally transformed cross-modal content generation. Pioneering architectures like LLaVA Liu et al. (2023) and Flamingo Alayrac et al. (2022) establish joint embedding spaces through cross-modal attention mechanisms, enabling unprecedented visual-linguistic synergy. These models achieve state-of-the-art performance in vision-language tasks ranging from contextual image captioning to visual commonsense reasoning, with recent extensions like Mini-Gemini Li et al. (2024b) demonstrating human-level multimodal comprehension. Liu and Bu (2024); Yoo et al. (2024) However, their rising capability to generate semantically coherent cross-modal content urgently demands robust solutions for intellectual property protection and content authenticity.
@@
We resolve these challenges through VLA-Mark, the first vision-language aligned watermarking framework that achieves cross-modally coordinated, quality-preserving watermark with excellent detectability and robustness via three innovations. First, extending beyond random vocabulary splitting, our Multiscale Semantic Saliency Metrics leverage visual semantics to guide green list selection through localized patch affinity (LPA), global semantic coherence (GSC), and cross-modal contextual salience (CCS). This aligns token partitioning with image content while maintaining zero training overhead. Second, our Entropy-Regulated Partition dynamically adjusts watermark intensity based on generation uncertainty and token criticality scores, prioritizing semantic preservation in low-entropy phases while enhancing watermark strength during high-entropy generation. Third, we introduce SCT based Distribution Adjustment through vision-aligned token prioritization, where cross-modal embedding alignment and fused metrics establish hierarchical protection for Semantic Critical Tokens (SCTs) against textual perturbations.
@@ 2.1 Cross-Modal Aligned Embedding
$\mathbf{Z}_{v}$ $=\text{VisEnc}(X_{v})=[\mathbf{z}_{\text{cls}};\mathbf{z}_{1},...,\mathbf{z}_{P}],$ (1)
where $\mathbf{Z}_{v}\in\mathbb{R}^{(P+1)\times d_{v}}$ and $P$ indicates the total number of image patch tokens augmented with a global [CLS] token. The subsequent alignment phase employs a trainable projection module $f_{\theta}(\cdot)$ , implemented as MLP Liu et al. (2024a) or generation adaptor Chen et al. (2025), to bridge the dimensional gap between modalities:
@@
where $f_{\theta}$ denotes parametric transformation that enables cross-modal compatibility while retaining original information patterns, so we get $\mathbf{H}_{v}\in\mathbb{R}^{(P+1)\times d}$ LLMs (e.g., Vicuna Chiang et al. (2023)) first tokenize input text of length $S$ and then retrieve text embeddings $\mathbf{H}_{q}\in\mathbb{R}^{S\times d}$ for LLM inference by querying the pretrained token embedding table, commonly referred to as the Vocabulary $\mathcal{V}$. We construct an embedding matrix $\mathbf{H}_{L}$ by removing non-linguistic elements such as symbols and numbers from $\mathcal{V}$, where $L$ denotes the number of linguistic tokens in the vocabulary. Then we use $\mathbf{H}_{v}$ and $\mathbf{H}_{L}$ in the following modules to find the SCT to guided $\mathcal{V}$ partitioning for watermark.
@@ 2.4 Entropy-Regulated Partition
$\mathcal{G}_{t}^{\text{SCT}}$ =$=\{w^{(1)},...,w^{(\lfloor\eta_{t}L\rfloor)}\},\quad$ (13)
@@ 2.5 SCT based Distribution Adjustment
where $p_{t}^{(k)}$ denotes the original logit value for token $k$ at step $t$, and $\delta>0$ controls the watermark intensity. This formulation applies: 1. Logit boosting (+$\delta$) for $\mathcal{G}_{t}$ tokens (SCT + green list) 2. Neutral treatment for $\mathcal{R}_{t}$ tokens (remaining vocabulary).
@@ 3.2.4 Text quality maintenance
In Figure 2 (left), we observe that our proposed approach exhibits lower median perplexity compared to other watermarking methods, indicating that it remains closer to the natural language distribution. This stems from our “semantic critical tokens,” which preserve core meanings and reduce unnecessary perturbations in high-salience tokens. In Figure 2 (right), average BLEU scores show that while all methods degrade as token length increases, our dynamic partitioning strategy and SCT protection help maintain relatively higher BLEU. By boosting tokens critical to the overall semantics, we minimize the distortion of fluency and coherence, leading to more faithful long generations.
@@ 3.3 Attack
Figure 3: AUC matrix for six watermarking methods under various attacks scenarios, with AUC values in parentheses. The proposed VLA-M retains high detection performance even under heavy text transformations such as paraphrasing and translation.
@@ 4.1 Text Watermarking Fundamentals
Contemporary watermarking techniques predominantly focus on unimodal text generation. The pioneering "green list" paradigm Kirchenbauer et al. (2023) partitions vocabulary through hash-based promotion, while entropy-aware variants Mao et al. (2024) modulate injection strength probabilistically. Distribution-preserving approaches Wu et al. (2024) maintain statistical fidelity through reweighting yet neglect semantic grounding. However, such unimodal designs fundamentally conflict with vision-conditioned generation: random vocabulary partitioning disrupts visual-semantic alignment by suppressing image-grounded tokens He et al. (2024), while static allocation strategies Liang et al. (2024) fail to adapt to cross-modal entropy variations Huang et al. (2023). Recent benchmarks Qiu et al. (2024) reveal 41% robustness degradation when deploying these methods in multimodal contexts, underscoring the necessity for vision-aligned watermark formulation.
@@ 4.2 Robustness Challenges and Attacks
Emerging adversarial attacks expose vulnerabilities through multimodal exploitation. Rastogi and Pruthi (2024) demonstrates 63% efficacy gain via black-box analysis-driven paraphrases, while He et al. (2024) reveals cross-lingual leakage during translation. Frameworks like DE-MARK Chen et al. (2024) remove watermarks via probabilistic n-gram erasure. Existing defenses remain unimodally confined—semantic preservation Ren et al. (2023) enhances robustness but cannot counter cross-modal attacks that jointly manipulate vision-text interdependencies. Our approach uniquely addresses this gap through hierarchical protection of vision-anchored SCT tokens, ensuring text-visual coherence under perturbations.
@@ 4.3 Vision-Language Aligned Architectures
State-of-the-art VLAMMs like LLaVA Liu et al. (2023) and BLIP-2 Li et al. (2023) establish cross-modal fusion through architectural innovations—gated cross-attention in Flamingo Alayrac et al. (2022) enables visual reasoning, while CogVLM2 Hong et al. (2024) leverages temporal grounding for scene understanding. Yet these models lack native authentication mechanisms, rendering generated content susceptible to adversarial attacks Rastogi and Pruthi (2024). Recent efforts Yoo et al. (2024) incorporate entropy adaptation but neglect alignment layers critical for coordinated embedding. Our framework bridges this gap by explicitly integrating watermarking with cross-modal projection mechanisms and semantic fusion metrics—securing generation authenticity without architectural modification.
@@ C.5 Robustness to Token Editing
$\Rightarrow\quad K$ $>N\left(1-\sqrt[-1]{(1-\gamma)}\right)$
@@
Time (seconds) VLA- Mark Cross Modal Aligned Embedding Multiscale Semantic Saliency Metrics Entropy Regulated Partition Fused Metric Guided Vocabulary SCT Distribution Adjustment All Components w/o watermark
@@ C.6 Visual-Semantic Coupling
(rank $\mathbb{P}(\text{rank change})\leq C\|\Delta X_{v}\|_{F}$ (59)
@@
(rank $\mathbb{P}(\text{rank preserved})\geq 1-C\|\Delta X_{v}\|_{F}\quad$ (60)

### v2

@@ Abstract
Vision-language models demand watermarking solutions that protect intellectual property without compromising multimodal coherence. Existing text watermarking methods disrupt visual-textual alignment through biased token selection and static strategies, leaving semantic-critical concepts vulnerable. We propose VLA-Mark, a vision-aligned framework that embeds detectable watermarks while preserving semantic fidelity through cross-modal coordination. Our approach integrates multiscale visual-textual alignment metrics, combining localized patch affinity, global semantic coherence, and contextual attention patterns, to guide watermark injection without model retraining. An entropy-sensitive mechanism dynamically balances watermark strength and semantic preservation, prioritizing visual grounding during low-uncertainty generation phases. Experiments show 7.4% lower PPL and 26.6% higher BLEU than conventional methods, with near-perfect detection (98.8% AUC). The framework demonstrates 96.1% attack resilience against attacks such as paraphrasing and synonym substitution, while maintaining text-visual consistency, establishing new standards for quality-preserving multimodal watermarking  [^1] Code is available at https://github.com/shiningwhite-cmd/VLA-mark.
@@ 1 Introduction
The emergence of vision-language aligned multimodal large models (VLAMMs) has fundamentally transformed cross-modal content generation. Pioneering architectures like LLaVA Liu et al. (2023) and Flamingo Alayrac et al. (2022) establish joint embedding spaces through cross-modal attention mechanisms, enabling unprecedented visual-linguistic synergy. These models achieve state-of-the-art performance in vision-language tasks ranging from contextual image captioning to visual commonsense reasoning Li et al. (2025), with recent extensions like Mini-Gemini Li et al. (2024b) demonstrating human-level multimodal comprehension. Liu and Bu (2024); Yoo et al. (2024); Ling et al. (2025) However, their rising capability to generate semantically coherent cross-modal content urgently demands robust solutions for intellectual property protection and content authenticity.
@@
We resolve these challenges through VLA-Mark, the first vision-language aligned watermarking framework that achieves cross-modally coordinated, quality-preserving watermark with excellent detectability and robustness via three innovations. First, extending beyond random vocabulary splitting, our Multiscale Semantic Saliency Metrics leverage visual semantics to guide green list selection through localized patch affinity (LPA), global semantic coherence (GSC) Hu et al. (2022), and cross-modal contextual salience (CCS). This aligns token partitioning with image content while maintaining zero training overhead. Second, our Entropy-Regulated Partition dynamically adjusts watermark intensity based on generation uncertainty and token criticality scores, prioritizing semantic preservation in low-entropy phases while enhancing watermark strength during high-entropy generation. Third, we introduce SCT based Distribution Adjustment through vision-aligned token prioritization, where cross-modal embedding alignment and fused metrics establish hierarchical protection for Semantic Critical Tokens (SCTs) against textual perturbations.
@@ 2.1 Cross-Modal Aligned Embedding
$\mathbf{Z}_{v}$ $=\text{VisEnc}(X_{v})=[\mathbf{z}_{\text{cls}};\mathbf{z}_{1},.,\mathbf{z}_{P}],$ (1)
where $\mathbf{Z}_{v}\in\mathbb{R}^{(P+1)\times d_{v}}$ and $P$ indicates the total number of image patch tokens augmented with a global [CLS] token. The subsequent alignment phase employs a trainable projection module $f_{\theta}(\cdot)$ , implemented as MLP Liu et al. (2024a) or generation adaptor Chen et al. (2025b), to bridge the dimensional gap between modalities:
@@
where $f_{\theta}$ denotes parametric transformation that enables cross-modal compatibility while retaining original information patterns, so we get $\mathbf{H}_{v}\in\mathbb{R}^{(P+1)\times d}$ . LLMs (e.g., Vicuna Chiang et al. (2023)) first tokenize input text of length $S$ and then retrieve text embeddings $\mathbf{H}_{q}\in\mathbb{R}^{S\times d}$ for LLM inference by querying the pretrained token embedding table, commonly referred to as the Vocabulary $\mathcal{V}$. We construct an embedding matrix $\mathbf{H}_{L}$ by removing non-linguistic elements such as symbols and numbers from $\mathcal{V}$, where $L$ denotes the number of linguistic tokens in the vocabulary. Then we use $\mathbf{H}_{v}$ and $\mathbf{H}_{L}$ in the following modules to find the SCT to guided $\mathcal{V}$ partitioning for watermark.
@@ 2.4 Entropy-Regulated Partition
$\mathcal{G}_{t}^{\text{SCT}}$ $=\{w^{(1)},...,w^{(\lfloor\eta_{t}L\rfloor)}\},\quad$ (13)
@@ 2.5 SCT based Distribution Adjustment
where $p_{t}^{(k)}$ denotes the original logit value for token $k$ at step $t$, and $\delta>0$ controls the watermark intensity. This formulation applies: 1. Logit boosting ($\delta$) for $\mathcal{G}_{t}$ tokens (SCT + green list) 2. Neutral treatment for $\mathcal{R}_{t}$ tokens (remaining vocabulary).
@@ 3.2.4 Text quality maintenance
In Figure 2 (left), we observe that our proposed approach exhibits lower median perplexity compared to other watermarking methods, indicating that it remains closer to the natural language distribution. This stems from our “semantic critical tokens,” which preserve core meanings and reduce unnecessary perturbations in high‐salience tokens. In Figure 2 (right), average BLEU scores show that while all methods degrade as token length increases, our dynamic partitioning strategy and SCT protection help maintain relatively higher BLEU. By boosting tokens critical to the overall semantics, we minimize the distortion of fluency and coherence, leading to more faithful long generations.
@@ 3.3 Attack
Figure 3: AUC matrix for six watermarking methods under various attacks scenarios, with AUC values in parentheses. The proposed VLA‐M retains high detection performance even under heavy text transformations such as paraphrasing and translation.
@@ 4.1 Text Watermarking Fundamentals
Contemporary watermarking techniques predominantly focus on unimodal text generation. The pioneering "green list" paradigm Kirchenbauer et al. (2023) partitions vocabulary through hash-based promotion, while other methods focus on sentence-level cohesion Zhang et al. (2025b), and entropy-aware variants Mao et al. (2024) modulate injection strength probabilistically. Distribution-preserving approaches Wu et al. (2024) maintain statistical fidelity through reweighting yet neglect semantic grounding. However, while watermarking is being explored for other modalities like video Huang et al. (2025), such unimodal designs fundamentally conflict with vision-conditioned generation: random vocabulary partitioning disrupts visual-semantic alignment by suppressing image-grounded tokens He et al. (2024), while static allocation strategies Liang et al. (2024) fail to adapt to cross-modal entropy variations Huang et al. (2023). Recent benchmarks and toolkits Qiu et al. (2024); Pan et al. (2024) reveal 41% robustness degradation when deploying these methods in multimodal contexts, underscoring the necessity for vision-aligned watermark formulation.
@@ 4.2 Robustness Challenges and Attacks
Emerging adversarial attacks expose vulnerabilities through multimodal exploitation. Rastogi and Pruthi (2024) demonstrates 63% efficacy gain via black-box analysis-driven paraphrases, while He et al. (2024) reveals cross-lingual leakage during translation. Frameworks like DE-MARK Chen et al. (2024) remove watermarks via probabilistic n-gram erasure. Existing defenses remain unimodally confined—semantic preservation Ren et al. (2023) enhances robustness, and related safety measures involve unlearning Chen et al. (2025a), but cannot counter cross-modal attacks that jointly manipulate vision-text interdependencies, a challenge highlighted in recent surveys Liu et al. (2025). Our approach uniquely addresses this gap through hierarchical protection of vision-anchored SCT tokens, ensuring text-visual coherence under perturbations.
@@ 4.3 Vision-Language Aligned Architectures
State-of-the-art VLAMMs, which represent a significant paradigm shift compared to earlier bidirectional models Zhang et al. (2025a), like LLaVA Liu et al. (2023) and BLIP-2 Li et al. (2023) establish cross-modal fusion through architectural innovations—gated cross-attention in Flamingo Alayrac et al. (2022) enables visual reasoning, while CogVLM2 Hong et al. (2024) leverages temporal grounding for scene understanding, with related work extending multimodality to speech Hei et al. (2025). Yet these models lack native authentication mechanisms, rendering generated content susceptible to adversarial attacks Rastogi and Pruthi (2024). Recent efforts Yoo et al. (2024) incorporate entropy adaptation but neglect alignment layers critical for coordinated embedding. Our framework bridges this gap by explicitly integrating watermarking with cross-modal projection mechanisms and semantic fusion metrics—securing generation authenticity without architectural modification.
@@ 5 Conclusion
Acknowledgement
We thank the reviewers for their valuable comments. This work was supported by the National Natural Science Foundation of China (Grant No.62506318); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.
@@ C.5 Robustness to Token Editing
[…]
@@
Time
VLA-
Mark
Cross
Modal
Aligned
Embedding
Multiscale
Semantic
Saliency
Metrics
Entropy
Regulated
Partition
Fused
Metric
Guided
Vocabulary
SCT
Distribution
Adjustment
All
Components
w/o
watermark
@@ C.6 Visual-Semantic Coupling
$\mathbb{P}(\text{rank change})\leq C\|\Delta X_{v}\|_{F}$ (59)
@@
$\mathbb{P}(\text{rank preserved})\geq 1-C\|\Delta X_{v}\|_{F}\quad$ (60)